# spectrum and local weak convergence of sparse random uniform hypergraphs

### v1

@@ 4.1.2. An exploration process in a rooted hypergraph
Let $H_{n}\in H(n,k,p)$. By 4.5, it is enough to know the distribution of $(G_{1}(H_{n}),1)$. We introduce a breadth-first search approach for a finite rooted hypergraph $(H_{n},1)$ to analyze the deviation of $(G_{1}(H_{n}),1)$ from the structure of $(GW_{k-1},o)$. This iterative procedure constructs a bijection $\phi:\mathcal{S}\to V(H_{n}(1))$, where $\mathcal{S}\subset\mathbb{N}_{k}^{\infty}$ and $V(H_{n}(1))$ denotes the vertex set of the connected component of $H_{n}$ containing the root vertex $i$. Given a vertex $j\in V(H_{n}(1))$, the pre-image $\phi^{-1}{(j)}$ is defined means the vertex $j$ is explored. The bijection $\phi$ is defined iteratively based on the ordering of $\mathbb{N}_{k}^{\infty}$ described earlier. We start with setting $\phi(o)=1$, the root, $A_{0}=\{1\}$, $C_{0}=\emptyset$, $U_{0}=V(H)\setminus\{1\}$, $CE_{0}=\emptyset$, $UE_{0}=E(H)$ and proceed in the following iteration:
We define $v_{t+1}=\phi(\mathbf{i}_{t})$, where $\mathbf{i}_{t}=\min\{\phi^{-1}(v):v\in A_{t}\}$, the minimum is with respect to the ordering on $\mathbb{N}^{\infty}_{\binom{n}{r}}$. The collection of unexplored hyperedges containing $v_{t+1}$ is denoted by $R_{t+1}=E_{v_{t+1}}(H)\cap UE_{t}$. Suppose that $\beta(t+1)=|R_{t+1}|$. If $\beta(t+1)>0$, we enumerate $R_{t+1}$ as
(15) $R_{t+1}=\{e_{t+1,1},e_{t+1,2},\ldots,e_{t+1,\beta(t+1)}\}.$
The set of unexplored neighbours of the vertex $v_{t+1}$ is denoted by
$I_{t+1}=\left(\bigcup\limits_{e\in R_{t+1}}e\right)\cap U_{t}.$
Note that, for two distinct $t_{1},t_{2}\geq 1$, without loss of generality suppose that $t_{1}<t_{2}$, we have $R_{t_{1}}\cap UE_{t_{1}}=\emptyset$ and $UE_{t_{2}}\subseteq UE_{t_{1}}$. Therefore the sets $R_{t_{1}}$ and $R_{t_{2}}$ are mutually disjoint set. Thus, if $a=|V(H)(1)|$ is the number of vertices in the connected component containing vertex $1$ then
$E(H(1))\subseteq\bigsqcup\limits_{t=1}^{b}R_{t}$
for some number $b\leq a$. Enumerating each $R_{t}$ induces an ordering on the edges of $H(1)$: for any two distinct edges $e_{t_{1},j},e_{t_{2},j^{\prime}}\in E(H(1))$, we write $e_{t_{1},j^{\prime}}<e_{t_{2},j}$ if
either (i) $t_{1}<t_{2}$, or (ii) $t_{1}=t_{2}$ and $j^{\prime}<j$.
We set $X_{t+1}=|I_{t+1}|$. In that case
(16) $X_{t+1}\leq(k-1)|R_{t+1}|.$
If $X_{t+1}>0$ then for any $u\in I_{t+1}$, suppose that $k=\min\{j:u\in e_{t+1,j}\in R_{t+1}\}$. We enumerate $U_{t}\cap e_{t+1,k}=\{u_{1},u_{2},\ldots\}$, and if $u=u_{r}$, then we define $\phi(\mathbf{i}_{t},(k,r))=u$, see Figure 4.
[…]
Figure 4. Iterative construction of the map $\phi$.
Now we set: for $t\in\mathbb{N}\cup\{0\}$,
@@
The knowledge gained at breadth-first exploration leads us to the filtration at $t$-th time step is represented by the filtration
@@
The iteration ended when $A_{t}=\emptyset$. Let $t_{0}$ denote the time when the process terminated, that is, $A_{t_{0}}=\emptyset$ but $A_{t_{0}-1}\neq\emptyset$. Then the vertex set $V(H)$ can be written as $V(H)=A_{t}\cup C_{t}\cup U_{t}$ for any $t\leq t_{0}$. Therefore, if $H$ is a hypergraph on $n$ vertex,
(18) $n=|A_{t}|+t+|U_{t}|.$
@@
Note that $d$-block Galton -Watson process the children of any individual never belongs to the past generations. Therefore if $(H_{n},1)\cong(T_{1},o)$ then, for any vertex $v_{t}$, each offspring hyperedge $e$ of $v_{t}$ satisfy the condition $e\setminus\{v_{t}\}\subseteq U_{t}$. Therefore, we always have $Y_{t}=Z_{t}$. Thus, the difference $Y_{t}-Z_{t}$ is one of the indications of deviation of $(H_{n},1)$ from $(T_{1},o)$. The following remarks give a road map of the proof of 4.4.
@@
    The difference between the degree distribution in the random hypergraph $H_{n}\in H(n,k,p)$ and the Poisson distribution with mean $\lambda$. Moreover, in a Galton-Watson $k$-hypertree, the degrees of different vertices are independent random variables. In contrast, in the random hypergraph $H_{n}$, the degree distributions of distinct vertices exhibit correlation. This discrimination is quantified in 4.9.
@@
      If $e$ and $e^{\prime}$ are not offspring of same vertex then $e\cap e^{\prime}=\{v\}$ implies that one of $e$ and $e^{\prime}$ is offspring hyperedge of $v$ and the other is the parent hyperedge of $v$.
    For example, the hypergraph illustrated in the Figure 5(b) is not a $3$-hypertree because if one starts exploring it from vertex $1$, then at the vertex $3$ we obtained two offspring hyperedges $\{3,5,6\}$ and $\{3,6,7\}$ of the vertex $3$ intersect at $2$ vertices. Furthermore, at vertex $5$, it is observed that $5\in\{3,5,6\}\cap\{5,6,8\}$ with $\{5,6,8\}$ is an offspring hyperedge of $5$ but the hyperedge $\{3,5,6\}$ is not a parent hyperedge of $5$. This contrast related to cyclic structure is formally captured in 4.10 and 4.11.
@@
(b) One offspring hyperedge of 33 intersect with the hyperedge $\{5,6,8\}$, which is offspring hyperedge of $3$. A pair of offspring hyperedges $\{3,5,6\}$ and $\{3,6,7\}$ of 33 intersecting in 22 vertices.
@@
Recall $D_{i}$ denotes the degree of the $i$-th vertex in the random uniform hypergraphs $H_{n}\in H(n,k,p)$. Observe that, from the discussion in Section 2.1.5, $D_{i}\sim\text{Bionomial}(r,p)$, where $r=\binom{n-1}{k-1}$. Therefore (20) leads us to the following equation:
@@
The equality $\{Z_{t}=Y_{t}\}$ of random variables represents the event in which, except one hyperedge, the vertex $v_{t+1}$ forms no other hyperedge with vertices outside of $U_{t}$. Therefore, by using 4.7, we have
@@
We write $a_{n}\approx b_{n}$ for large $n$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$. As for large $n$, it holds that $(\binom{n-1}{k-1}-\binom{|U_{t}|-1}{k-1})\approx\left((n-1)^{k-1}-(|U_{t}|-1)^{k-1}\right)\frac{1}{(k-1)!}$. Therefore, for large $n$, using 4.8 we have the following:
@@
Again for large $n$, we have $\left(\binom{n-1}{k-1}-\binom{|U_{t}|}{k-1}\right)\approx((n-1)^{k-1}-(|U_{t}|)^{k-1})\frac{1}{(k-1)!}$. By the similar calculation as in the proof of 4.9, for some $c_{1}>0$, we have
@@
Again, as shown in the proof of 4.9, we have $\mbox{\bf E}(|A_{t}|)\leq 1+(\lambda-1)t$. Therefore
@@
For large $n$, using 4.8 we have
@@
Now we are ready to give a proof of 4.1 using 4.9, 4.10 and 4.11.
Proof of 4.4.
@@
$E_{1}$ $=\{(Z^{\prime}_{1},\ldots,Z^{\prime}_{t_{0}\wedge\tau})\neq(Y_{1},\ldots,Y_{t_{0}\wedge\tau})\}.$
$E_{2}$ $=\{\exists 0\leq s\leq t_{0}\wedge\tau\;:\;|e\backslash U_{t}|>1\mbox{ for some }e\in R_{t}\}.$
$E_{3}$ $=\{\exists 0\leq s\leq t_{0}\wedge\tau\;:\;|e\cap e^{\prime}|>1\mbox{ for some }e\neq e^{\prime}\in R_{t}\}.$
As discussed in 4.6, the deviation described by (24) depends on the following two reasons: Firstly, observe that ${\bf P}(G_{1}(GW_{k-1},o)_{\lceil\frac{1}{\delta}\rceil}\cong g_{\lceil\frac{1}{\delta}\rceil})\neq 0$ only if $g=G_{1}(T)$ for some (k1)(k-1)-block tree $T$ of depth $t_{0}$ from its root. The difference in (24) is non-zero because of the occurrence of event $E_{1}$.
Secondly, for d=k1d=k-1, the $d$-block Galton-Watson tree $(GW_{d},o)$ admits a natural representation of a $1$-set line graph of an $1$-set rooted $k$-uniform hypergraph $(T_{k},o)$. The main characteristics of $(T_{k},o)$ is for two hyperedges $e$ and $e^{\prime}$, we must have $|e\cap e^{\prime}|\leq 1$. The $|e\cap e^{\prime}|=1$ occurs with the following two cases: (1)if $e$ and $e^{\prime}$ are two offspring hyperedges of a vertex $v$ then $e\cap e^{\prime}=\{v\}$, and (2) if $e$ and $e^{\prime}$ are not the offspring of the same vertex then $e\cap e^{\prime}=\{v\}$ implies one of $e$ and $e^{\prime}$ is offspring hyperedge of $v$ and the other is the parent hyperedge of $e$. The deviation of (24) also depends on the fact how much $(G_{1}(H_{n}(1)),i)_{l}$ deviates from this characteristics of $(T_{k},o)$. In other words, the difference in (24) is non-zero if $E_{2}\cup E_{3}$ occurs. Thus, we have
$\left|\mu_{n}(B_{\delta}(g))-\nu_{k-1,\lambda}(B_{\delta}(g))\right|\leq{\bf P}(E_{1})+{\bf P}(E_{2})+{\bf P}(E_{3}).$
Again, by the union bounds from 4.10 and 4.11 we have
${\bf P}(E_{2})\leq\frac{c_{1}(t_{0}\wedge\tau)^{2}}{n-1}\mbox{ and }{\bf P}(E_{3})$ $\leq\frac{c_{2}(t_{0}\wedge\tau)}{n-1}.$
Therefore, using 4.10 and $t_{0}\wedge\tau\leq t_{0}$, we get
@@ 4.2. Almost sure convergence
This subsection is dedicated to prove Section 4.2. The proof (which is partially motivated by [23, Lemma 2.3 and Lemma 2.4] and [52, Section 8.4]), is devided in two steps. In the first step, we show that for large enough $n$, under the 1, the number of hyperedges in $H_{n}\in H(n,k,p)$ is bounded by $\frac{3n\lambda}{2k}$ almost surely. See 4.12. Using this fact, in the second step, we construct a martingale sequence. That martingale sequence leads us to our desired result using Azuma Inequality [21, Theorem 5.2]. See 4.13.
@@
Let $\lambda,k,r,n,p$ be as in Assumption 1, and $\xi_{n}$ be as defined in (25). Then there exists a positive constant $c_{4}$ such that
@@
Next, we consider the lexicographic ordering of the set $\left[\begin{subarray}{c}n\\ k\end{subarray}\right]=\{e_{1},\ldots,e_{r}\}$, where $r=\binom{n}{k}$. We define the following random variables:
@@
Note that $\mathcal{E}_{j}$ reveals the information of the first $j$ hyperedges of $H_{n}\in H(n,k,p)$ in lexicographic ordering. Observe that, the sigma algebra $\mathcal{E}_{j}$ contains the information about both the hyperedges and non-hyperedges in $H_{n}\in H(n,k,p)$ up to $e_{E_{j}}$.
@@
$\Big{|}\mbox{\bf E}[S_{n}|\mathcal{E}_{i+1}]-\mbox{\bf E}[S_{n}|\mathcal{E}_{i}]\Big{|}\leq\frac{2k}{n}\sum\limits_{j=1}^{N}\lambda^{i},\text{ ~almost surely}.$
Now we recall a few well known results, such as, the Doob’s Martingale Property, the Azuma-Hoeffding inequality and the Borel-Cantelli lemma, which will be used in the proof of Section 4.2. Let $(\Omega,\mathcal{F},{\bf P})$ be a probability space, and let $\{\mathcal{F}_{n}\}_{n\geq 0}$ be a filtration, i.e., an increasing sequence of sigma-algebras:
@@
We proceed to prove Section 4.2 assuming 4.12 and 4.13. The proofs of these lemmas will be given at the end of this section.
Proof of Section 4.2.
@@
Again, for any $A\in\mathcal{B}_{m}$, we have
@@
It clear that $\mbox{\bf E}[|S_{n}|]\leq 1$. Therefore the Doob’s Martingale Property, 1, implies that $\{M_{i}\}$ is a martingale with respect to the filtration $\{\mathcal{E}_{i}\}$.
Let $c=2k\sum\limits_{j=1}^{m}\lambda^{i}$. Then from 4.13 we get, for $i=0,1,\ldots,m$,
@@
Therefore 2 implies that, for $N=\frac{3n\lambda}{2k}$,
@@
where $c_{3}=\epsilon^{2}k/3\lambda c^{2}$. Note that if $\{Z_{n}\leq N\}$ holds then $S_{n}$ is $\mathcal{E}_{N}$ measurable. Therefore, if $\{Z_{n}\leq N\}$ occurs then $M_{N}=S_{n}$. Also we have $\mbox{\bf E}[M_{N}]=\mbox{\bf E}[S_{n}]$. Therefore, by 4.12 and (27), we get
@@
Then Borel-Cantelli Lemma, 3, implies that
@@
for all $A\in\mathcal{B}_{m}$. The Portmanteau theorem (4.3) implies that
@@
It remains to prove 4.12 and 4.13. We recall the Chernoff Bound, which will be used in the proof of 4.12.
@@
Proof of 4.12.
Note that $\xi_{n}$ is a sum of independent Bernoulli random variables. Then using the Chernoff Bound, 4, we have
@@
Again, we have $\mbox{\bf E}[\xi_{n}]=\binom{n}{k}p=n\lambda/k$. The result follows by putting $\delta=1/2$. ∎
Proof of 4.13.
Note that we have
@@
Let $N=\frac{3\lambda n}{2k}$. 4.12 and 3 implies that the number of hyperedges in $H_{n}$ is at most $N$ almost surely, for large $n$. Thus, if the sigma-algebra $\mathcal{E}_{i}$ exposes the $i$-th hyperedge $e$, then the root vertex belongs to one of the $N$-radius balls centred at some vertex belonging to $e$ (see Figure 6).
@@
choices for being a root which is affected by revealing the $i$-th hyperedge. Since for all $j=1,\ldots,n$, the difference
@@
Recall 1. If we relax the condition $\binom{n-1}{k-1}p=\lambda$ to $\binom{n-1}{k-1}p\to\lambda,$ the same argument remains valid. In this regime, the binomial distribution $\mathrm{Bin}(\binom{n-1}{k-1},p)$ converges in law to $\mathrm{Poi}(\lambda)$, so the limiting distribution is again $\Gamma_{k,\lambda}$. Furthermore, the concentration inequalities continue to hold uniformly as long as $\binom{n-1}{k-1}p$ remains bounded, which also yields almost sure convergence. Hence, 3.1 is preserved under the weaker assumption $\binom{n-1}{k-1}p\to\lambda$. Consequently, 3.2 follows directly from 3.1.

### v2

@@ 4.1.2. An exploration process in a rooted hypergraph
Let $H_{n}\in H(n,k,p)$. By Lemma 4.5, it is enough to know the distribution of $(G_{1}(H_{n}),1)$. We introduce a breadth-first search approach for a finite rooted hypergraph $(H_{n},1)$ to analyze the deviation of $(G_{1}(H_{n}),1)$ from the structure of $(GW_{k-1},o)$. This iterative procedure constructs a bijection $\phi:\mathcal{S}\to V(H_{n}(1))$, where $\mathcal{S}\subset\mathbb{N}_{k}^{\infty}$ and $V(H_{n}(1))$ denotes the vertex set of the connected component of $H_{n}$ containing the root vertex $i$. Given a vertex $j\in V(H_{n}(1))$, we declare $j$ explored once $\phi^{-1}(j)$ is defined. The bijection $\phi$ is defined iteratively based on the ordering of $\mathbb{N}_{k}^{\infty}$ described earlier. We start by setting $\phi(o)=1$, the root; $A_{0}=\{1\}$, the initial active set; $C_{0}=\emptyset$, the initial covered component; $U_{0}=V(H)\setminus\{1\}$, the initial unexplored component; $CE_{0}=\emptyset$, the initial covered hyperedge set; $UE_{0}=E(H)$, the initial unexplored hyperedge set; and proceed in the following iteration:
@@
Where $I_{t}$ and $R_{t}$ are defined in the following discussion. We define $v_{t+1}=\phi(\mathbf{i}_{t})$, where $\mathbf{i}_{t}=\min\{\phi^{-1}(v):v\in A_{t}\}$, the minimum is with respect to the ordering on $\mathbb{N}^{\infty}_{\binom{n}{r}}$. The collection of unexplored hyperedges containing $v_{t+1}$ is denoted by
$R_{t+1}=E_{v_{t+1}}(H)\cap UE_{t}.$
Suppose that $\beta(t+1)=|R_{t+1}|$. If $\beta(t+1)>0$, we enumerate $R_{t+1}$ as
(15) $R_{t+1}=\{e_{t+1,1},e_{t+1,2},\ldots,e_{t+1,\beta(t+1)}\}.$
The set of unexplored neighbours of the vertex $v_{t+1}$ is given by
(16) $I_{t+1}=\left(\bigcup\limits_{e\in R_{t+1}}e\right)\cap U_{t}.$
We set $X_{t+1}=|I_{t+1}|$ for $t\in\mathbb{N}\cup\{0\}$. In that case by Equation 16
(17) $X_{t+1}\leq(k-1)|R_{t+1}|.$
Iteration of the bijection $\phi$
To describe the iteration of $\phi$, we need to enumerate the vertex set of $H_{n}$ using the breadth-first-search.
Note 4.6.
The hyperedge set of the connected component of $H_{n}$ that contains the root can be expressed as disjoint union of $R_{t}$.
Proof.
For two distinct $t_{1},t_{2}\geq 1$, without loss of generality suppose that $t_{1}<t_{2}$. In that case, we have $R_{t_{1}}\cap UE_{t_{1}}=\emptyset$ and $UE_{t_{2}}\subseteq UE_{t_{1}}$. Therefore the sets $R_{t_{1}}$ and $R_{t_{2}}$ are mutually disjoint set. Thus, if $a=|V(H)(1)|$ is the number of vertices in the connected component containing vertex $1$ then
$E(H(1))=\bigsqcup\limits_{t=1}^{b}R_{t}$
for some number $b\leq a$. ∎
[…]
Figure 4. Iterative construction of the map $\phi$.
Note 4.7.
Enumeration of $R_{t}$ provide us the iterative definition of $\phi$.
Proof.
Enumerating each $R_{t}$ induces an ordering on the edges of $H(1)$: for any two distinct edges $e_{t_{1},j},e_{t_{2},j^{\prime}}\in E(H(1))$, we write $e_{t_{1},j^{\prime}}<e_{t_{2},j}$ if
either (i) $t_{1}<t_{2}$, or (ii) $t_{1}=t_{2}$ and $j^{\prime}<j$.
If $X_{t+1}>0$ then for any $u\in I_{t+1}$, suppose that $k=\min\{j:u\in e_{t+1,j}\in R_{t+1}\}$. We enumerate $U_{t}\cap e_{t+1,k}=\{u_{1},u_{2},\ldots\}$, and if $u=u_{r}$, then we define $\phi(\mathbf{i}_{t},(k,r))=u$, see Figure 4. ∎
Let $t_{0}$ denote the time when the process terminated, that is, $A_{t_{0}}=\emptyset$ but $A_{t_{0}-1}\neq\emptyset$. For $t<t_{0}$, the knowledge gained at breadth-first exploration leads us to the filtration at $t$-th time step is represented by the filtration
@@
The iteration ended when $A_{t}=\emptyset$. Then the vertex set $V(H)$ can be written as $V(H)=A_{t}\cup C_{t}\cup U_{t}$ for any $t\leq t_{0}$. By definition of the iteration, exactly one new vertex is added to $C_{t}$ at each step. Thus, $|C_{t}|=t$. Therefore, if $H$ is a hypergraph on $n$ vertex, $n=|A_{t}|+t+|U_{t}|$. ∎
@@
Note that $d$-block Galton -Watson process the children of any individual never belongs to the past generations. Therefore if $(H_{n},1)\cong(T_{1},o)$ then, for any vertex $v_{t}$, each offspring hyperedge $e$ of $v_{t}$ satisfy the condition $e\setminus\{v_{t}\}\subseteq U_{t}$. Therefore, we always have $Y_{t}=Z_{t}$. Thus, the difference $Y_{t}-Z_{t}$ is one of the indications of deviation of $(H_{n},1)$ from $(T_{1},o)$. The following remarks give a road map of the proof of Proposition 4.4.
@@
    The difference between the degree distribution in the random hypergraph $H_{n}\in H(n,k,p)$ and the Poisson distribution with mean $\lambda$. Moreover, in a Galton-Watson $k$-hypertree, the degrees of different vertices are independent random variables. In contrast, in the random hypergraph $H_{n}$, the degree distributions of distinct vertices exhibit correlation. This discrimination is quantified in Lemma 4.12.
@@
      If $e$ and $e^{\prime}$ are not offspring of the same vertex, then $e\cap e^{\prime}=\{v\}$ implies that one of $e$ and $e^{\prime}$ is offspring hyperedge of $v$ and the other is the parent hyperedge of $v$.
    For example, the hypergraph illustrated in the Figure 5(b), if one starts exploring it from vertex $1$, then at the vertex $3$ we obtained two offspring hyperedges $\{3,5,6\}$ and $\{3,6,7\}$ of the vertex $3$ intersect at $2$ vertices. Furthermore, at vertex $5$, it is observed that $5\in\{3,5,6\}\cap\{5,6,8\}$ with $\{5,6,8\}$ is an offspring hyperedge of $5$ but the hyperedge $\{3,5,6\}$ is not a parent hyperedge of $5$. This contrast related to cyclic structure is formally captured in Lemma 4.13 and Lemma 4.14.
@@
(b) One offspring hyperedge of 33 intersect with the hyperedge $\{5,6,8\}$, which is offspring hyperedge of $5$. A pair of offspring hyperedges $\{3,5,6\}$ and $\{3,6,7\}$ of 33 intersecting in 22 vertices.
@@
Recall $D_{i}$ denotes the degree of the $i$-th vertex in the random uniform hypergraphs $H_{n}\in H(n,k,p)$. Observe that, from the discussion in Section 2.1.5, $D_{i}\sim\text{Binomial}(r,p)$, where $r=\binom{n-1}{k-1}$. Therefore (21) leads us to the following equation:
@@
The equality $\{Z_{t}=Y_{t}\}$ of random variables represents the event in which, except one hyperedge, the vertex $v_{t+1}$ forms no other hyperedge with vertices outside of $U_{t}$. Therefore, by using Lemma 4.10, we have
@@
We write $a_{n}\approx b_{n}$ for large $n$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$. As for large $n$, it holds that $(\binom{n-1}{k-1}-\binom{|U_{t}|-1}{k-1})\approx\left((n-1)^{k-1}-(|U_{t}|-1)^{k-1}\right)\frac{1}{(k-1)!}$. Therefore, for large $n$, using Lemma 4.11 we have the following:
@@
Again for large $n$, we have $\left(\binom{n-1}{k-1}-\binom{|U_{t}|}{k-1}\right)\approx((n-1)^{k-1}-(|U_{t}|)^{k-1})\frac{1}{(k-1)!}$. By the similar calculation as in the proof of Lemma 4.12, for some $c_{1}>0$, we have
@@
Again, as shown in the proof of Lemma 4.12, we have $\mbox{\bf E}(|A_{t}|)\leq 1+(\lambda-1)t$. Therefore
@@
For large $n$, using Lemma 4.11 we have
@@
Now we are ready to give a proof of Proposition 4.1 using Lemma 4.12, Lemma 4.13 and Lemma 4.14.
Proof of Proposition 4.4.
@@
$A_{1}$ $=\{(Z^{\prime}_{1},\ldots,Z^{\prime}_{t_{0}\wedge t})\neq(Y_{1},\ldots,Y_{t_{0}\wedge t})\}.$
$A_{2}$ $=\{\exists 0\leq s\leq t_{0}\wedge t\;:\;|e\backslash U_{t}|>1\mbox{ for some }e\in R_{t}\}.$
$A_{3}$ $=\{\exists 0\leq s\leq t_{0}\wedge t\;:\;|e\cap e^{\prime}|>1\mbox{ for some }e\neq e^{\prime}\in R_{t}\}.$
As discussed in Remark 4.9, the deviation described by (25) depends on the following two reasons: Firstly, observe that ${\bf P}(G_{1}(GW_{k-1},o)_{\lceil\frac{1}{\delta}\rceil}\cong g_{\lceil\frac{1}{\delta}\rceil})\neq 0$ only if $g=G_{1}(T)$ for some (k1)(k-1)-block tree $T$ of depth $t_{0}$ from its root. The difference in (25) is non-zero because of the occurrence of event $A_{1}$.
Secondly, for d=k1d=k-1, the $d$-block Galton-Watson tree $(GW_{d},o)$ admits a natural representation of a $1$-set line graph of an $1$-set rooted $k$-uniform hypergraph $(T_{k},o)$. The main characteristics of $(T_{k},o)$ is for two hyperedges $e$ and $e^{\prime}$, we must have $|e\cap e^{\prime}|\leq 1$. The $|e\cap e^{\prime}|=1$ occurs with the following two cases: (1)if $e$ and $e^{\prime}$ are two offspring hyperedges of a vertex $v$ then $e\cap e^{\prime}=\{v\}$, and (2) if $e$ and $e^{\prime}$ are not the offspring of the same vertex then $e\cap e^{\prime}=\{v\}$ implies one of $e$ and $e^{\prime}$ is offspring hyperedge of $v$ and the other is the parent hyperedge of $e$. The deviation of (25) also depends on the fact how much $(G_{1}(H_{n}(1)),i)_{l}$ deviates from this characteristics of $(T_{k},o)$. In other words, the difference in (25) is non-zero if $A_{2}\cup A_{3}$ occurs. Thus, we have
$\left|\mu_{n}(B_{\delta}(g))-\nu_{k-1,\lambda}(B_{\delta}(g))\right|\leq{\bf P}(A_{1})+{\bf P}(A_{2})+{\bf P}(A_{3}).$
Again, by the union bounds from Lemma 4.13 and Lemma 4.14 we have
${\bf P}(A_{2})\leq\frac{c_{1}(t_{0}\wedge t)^{2}}{n-1}\mbox{ and }{\bf P}(A_{3})$ $\leq\frac{c_{2}(t_{0}\wedge t)}{n-1}.$
Therefore, using Lemma 4.13 and $t_{0}\wedge t\leq t_{0}$, we get
@@ 4.2. Almost sure convergence
This subsection is dedicated to prove Proposition 4.2. The proof (which is partially motivated by [23, Lemma 2.3 and Lemma 2.4] and [55, Section 8.4]), is divided in two steps. In the first step, we show that for large enough $n$, under the Assumption 1, the number of hyperedges in $H_{n}\in H(n,k,p)$ is bounded by $\frac{3n\lambda}{2k}$ almost surely. See Lemma 4.15. Using this fact, in the second step, we construct a martingale sequence. This martingale leads to the desired result using Azuma Inequality [21, Theorem 5.2]. See Lemma 4.16.
@@
Let $\lambda,k,r,n,p$ be as in Assumption 1, and $\xi_{n}$ be as defined in (26). Then
@@
Next, we consider the lexicographic ordering of the set $\left[\binom{\begin{subarray}{c}n\end{subarray}}{k}\right]=\{e_{1},\ldots,e_{r}\}$, where $r=\binom{n}{k}$. We define the following random variables:
@@
Note that $\mathcal{E}_{j}$ reveals the information of the first $j$ hyperedges of $H_{n}\in H(n,k,p)$ in lexicographic ordering. Observe that, the sigma-algebra $\mathcal{E}_{j}$ contains the information about both the hyperedges and non-hyperedges in $H_{n}\in H(n,k,p)$ up to $e_{j}$.
@@
$\Big|\mbox{\bf E}[S_{n}|\mathcal{E}_{i+1}]-\mbox{\bf E}[S_{n}|\mathcal{E}_{i}]\Big|\leq\frac{2k}{n}\sum\limits_{j=1}^{N}\lambda^{j},\text{ ~almost surely}.$
Now we recall a few well known results, such as, the Doob’s Martingale Property, the Azuma-Hoeffding inequality and the Borel-Cantelli lemma, which will be used in the proof of Proposition 4.2. Let $(\Omega,\mathcal{F},{\bf P})$ be a probability space, and let $\{\mathcal{F}_{n}\}_{n\geq 0}$ be a filtration, i.e., an increasing sequence of sigma-algebras:
@@
We proceed to prove Proposition 4.2 assuming Lemma 4.15 and Lemma 4.16. The proofs of these lemmas will be given at the end of this section.
Proof of Proposition 4.2.
@@
Again, for any $A\in\mathcal{B}_{m}$, note that
@@
It is clear that $\mbox{\bf E}[|S_{n}|]\leq 1$. Therefore the Doob’s Martingale Property, Fact 1, implies that $\{M_{i}\}$ is a martingale with respect to the filtration $\{\mathcal{E}_{i}\}$.
Let $c=2k\sum\limits_{j=1}^{m}\lambda^{i}$. Then from Lemma 4.16 we get, for $i=0,1,\ldots,m$,
@@
Therefore Fact 2 implies that, for $N=\frac{3n\lambda}{2k}$,
@@
where $c_{3}=\epsilon^{2}k/3\lambda c^{2}$. Note that if $\{Z_{n}\leq N\}$ holds then $S_{n}$ is $\mathcal{E}_{N}$ measurable. Therefore, if $\{Z_{n}\leq N\}$ occurs then $M_{N}=S_{n}$. Also observe that $\mbox{\bf E}[M_{N}]=\mbox{\bf E}[S_{n}]$. Therefore, by Lemma 4.15 and (28), we get
@@
Then Borel-Cantelli Lemma, Fact 3, implies that
@@
for all $A\in\mathcal{B}_{m}$. The Portmanteau theorem (Theorem 4.3) implies that
@@
It remains to prove Lemma 4.15 and Lemma 4.16. We recall the Chernoff Bound, which will be used in the proof of Lemma 4.15.
@@
Proof of Lemma 4.15.
Note that $\xi_{n}$ is a sum of independent Bernoulli random variables. Then using the Chernoff Bound, Fact 4, we have
@@
Again, note that $\mbox{\bf E}[\xi_{n}]=\binom{n}{k}p=n\lambda/k$. The result follows by putting $\delta=1/2$. ∎
Proof of Lemma 4.16.
Note that
@@
Let $N=\frac{3\lambda n}{2k}$. Lemma 4.15 and Fact 3 implies that the number of hyperedges in $H_{n}$ is at most $N$ almost surely, for large $n$. Thus, if the sigma-algebra $\mathcal{E}_{i}$ exposes the $i$-th hyperedge $e$, then the root vertex belongs to one of the $N$-radius balls centred at some vertex belonging to $e$ (see Figure 6).
@@
choices for the root being affected by revealing the $i$-th hyperedge. Since for all $j=1,\ldots,n$, the difference
@@
Recall Assumption 1. If we relax the condition $\binom{n-1}{k-1}p=\lambda$ to $\binom{n-1}{k-1}p\to\lambda,$ the same argument remains valid. In this regime, the binomial distribution $\mathrm{Bin}(\binom{n-1}{k-1},p)$ converges in law to $\mathrm{Poi}(\lambda)$, so the limiting distribution is again $\Gamma_{k,\lambda}$. Furthermore, the concentration inequalities continue to hold uniformly as long as $\binom{n-1}{k-1}p$ remains bounded, which also yields almost sure convergence. Hence, Theorem 3.1 is preserved under the weaker assumption $\binom{n-1}{k-1}p\to\lambda$. Consequently, Corollary 3.2 follows directly from Theorem 3.1.